What does plus or minus a standard deviation of 17 cm mean?
It refers to some variable whose distribution has got some mean and the measure of its spread is a standradr deviation of 17 cm. If the distribution is Normal (or approximately so), then the interval from the mean measure minus 17 cm to the mean plus 17 cm will contain approx 68% of the total number of observations.
How do you do you increase accuracy in a theoritcal or experimental probability?
You improve your model through a better understanding of the underlying processes.
Although more trials will improve the accuracy of experimental probability they will make no difference to the theoretical probability.
There are five letters, and two of them are s's. The theoretical probability of choosing an s would be 2 out of 5.
2/5 or 40%
What is the probability of getting at least one head if two coins were flipped?
Let put the question in an other way : obtaining at least one head is the contrary of obtaining two tails at the same time.
The probability to obtain one tail with first coin is 1/2,
the probability to obtain one tail with the other is also 1/2,
so the probability to obtain one tail on each coin is 1/2x1/2 = 1/4
Thus the probability to obtain at least one head is 1-1/4=3/4
What if you lost your nso imo roll number?
there is only 1 way that you should ask your school principle
What is the probability of throwing 2 dice so that you have a total score of 5?
When you throw 2 6-sided dice, there are 36 ways they can come up. Of those, 4 (1 and 4, 2 and 3, 3 and 2, 4 and 1) total 5. So the probability is 4 chances in 36, or 1 in 9 (1/9 = 0.11111...)
When you toss a coin straight up the acceleration of the coin is?
As soon as you let go of the coin, there will be no upwards force acting on the coin and so no upward acceleration. There will be downwards acceleration due to the force of the earth's gravity.
If you want to be more accurate you could include aerodynamic drag which will operate in the direction opposite to that of the motion.
Pavlovian response.
How do simulations help you understand the probability of events happening?
by keeping in touch with your family and friends and helping others
What is the probability of getting two heads in a row followed by three tails in a row?
If you toss a coin often enough, it is a certainty, so that the probability = 1.
The probability of that outcome in the first five tosses of a coin is (1/2)5 = 1/32.
"Probability" =
the chance that an event either will or won't happen. Every event either
will or won't happen, so the sum of the two chances is ' 1 '.
"Odds" =
the ratio of the (probability that it will happen) to the (probability that it won't).
-- If (happening) and (not-happening) are equally likely, then each probability
is 0.5, and odds are 0.5/0.5 = 1 .
-- If (happening) is more likely than (not-happening) then probability of happening
is more than 0.5, and probability of not-happening is less than 0.5.
Their sum is still ' 1 ', because there is a 100% chance that the event will either happen
or not happen.
But the odds are now (more than 0.5)/(less than 0.5) = more than 1 .
What are the odds of getting 1 2 3 4 5 6 if you roll one six sided dice?
1/6 for all 6 numbers.
-- On the first roll, the probability of a ' 1 ' is 1/6 .
-- On the second roll, the probability of a ' 2 ' is 1/6 .
-- On the third roll, the probability of a ' 3 ' is 1/6 .
-- On the fourth roll, the probability of a ' 4 ' is 1/6 .
-- On the fifth roll, the probability of a ' 5 ' is 1/6 .
-- On the sixth roll, the probability of a ' 6 ' is 1/6 .
So the probability of rolling 1 - 2 - 3 - 4 - 5 - 6 in six rolls is (1/6)6 = 0.000021433 = 0.0021%
What is the probability of the first and second card an ace?
Assuming it is a standard poker deck with 52 cards and 4 aces
The probability of getting your first card an ace is = 4/52
Over here you need to reread your questions. There are something you need to know before i continue - when you draw your second card, did you return the first card back into the pile? lets say if you draw an ace of spades, would you be able to redraw it again for the second card?
If the first ace is return to the pile:
probability of getting a second ace is also 4/52
so the total probability of getting both cards an ace is (4/52 x 4/52)
If the first ace is not return to the pile:
probability of getting a second ace is now 3/51
note that removing one ace also removes one card from the pile. total card is now 51 with 3 aces only
so the total probability of getting both cards an ace is (4/52 x 3/51)
hope i help.
1/33000
What to do when a continuous probability is used to approximate a discrete probability distribution?
You may need to make a continuity correction - see statistical text books for details.
Half the discrete unit.
How do you apply standard deviation to normal curve?
The distance between the middle and the inflection point is the standard deviation.
What is the probability of picking an ace or picture card from a deck of cards?
If you pick 37 cards without replacement, or pick a card from the bottom of a mint deck, the probability is 1: it is a certainty.
If you pick a random card from a deck, then the probability is 4/13.
Does a negative z score always yield a negative percentile?
A percentile is always between 0% and 100%, never negative.
What are the probabilities of getting a sum o f7?
The experiment is undefined. On a single roll of eight or more normal dice the probability of getting a sum of 7 is 0.